Tuesday, January 31, 2006

Change 0.4uF cap for thyratron trigger

Keith suggested to decrease the voltage by 15 kV for passivation. I tried this way and found the system worked well, and the red light from the chamber looked homogeneous. However the south side randomly missed triggers, I can hear the click from the south thyratrons and see the sparks inside the thyratrons. I thought maybe the trigger circiut board send the noise, so I removed the circiut board. Casually I found the capacitor connected between thyraton CX1585 and CX1622 was very dirty, there was some yellow liqud leaked from the capacitor. I changed this dirty one by a clean one.

So far the passivation has been run well.

Monday, January 30, 2006

What caused the resistor burnt ?

Discussed with Mike the problem of resisor burning. We thought there were two parts to induce the big current pass through the 100k Ohm resistors. One is the 6 thyratrons, the other is the 5 caps.

I also 1) Replaced 10k Ohm resistor with 15k Ohm, and ordered 10k Ohme resistors. 2) Found 60A fuse blow off, and changed it.

Friday, January 27, 2006

The 10k Ohm resistor broken again

I tried to passivate the system today. Unfortunately, I found the connect wire was burnning after running half hour, this phenomenon is very similar to what happened on December 06, 2005.

Maybe we didn't find the main problem last time? What's caused the resistor burned?

Thursday, January 26, 2006

Passivation Prometheus system

The system has been run under 1.2 bar Helium gas for a copule of minutes. Then I filled a little Flourine gas for passivation. The passivation process has been kept for about more than one hour, the system ran very stable.

Science News: Ion beams feel the squeeze

Nature Editor's Summary: Ion beams feel the squeeze

Acceleration of particles by intense laser plasma interactions is a promising alternative to conventional particle accelerators. Two groups this week report advances in this field that bring the prospect of monoenergetic ion beams closer. Such beams are necessary for various potential applications including medical proton and heavy-ion therapy. Hegelich et al. produced laser-driven C5+ions with a vastly reduced energy spread compared to previous experiments. And Schwoerer et al. produced quasi-monoenergetic proton beams from intense laser irradiation of solid microstructured targets.

  1. B. M. Hegelich, B. J. Albright, J. Cobble, K. Flippo, S. Letzring, M. Paffett, H. Ruhl, J. Schreiber, R. K. Schulze and J. C. Fernández, Laser acceleration of quasi-monoenergetic MeV ion beams, Nature, vol. 439, 441(2006).
  2. H. Schwoerer, S. Pfotenhauer, O. Jackel, K.-U. Amthor, B. Liesfeld, W. Ziegler, R. Sauerbrey, K. W. D. Ledingham and T. Esirkepov, Laser?plasma acceleration of quasi-monoenergetic protons from microstructured targets, Nature, vol. 439, 445(2006).

Wednesday, January 25, 2006


Prometheus system running well!

I vacuumed the gas reservior and filled Helium gas by 1.2 bar. After warming up the system for 15 minutes, I turned on the power. The x-ray gun was triggered normally and the x-ray flux could be measured. The Anode voltgae monitor showed 150 kV. The rail gap triggered well, I could apply 30 kV to main gaps.

But I found a small problem, there were acrs in the north side, which produced carbon track in the top oil tank. I pumped the oil and found the connects not well. It's easy to fix. I pumped the oil out and double cleaned. It's almost 8:00pm after pumping oil back, I will try to run the system tomorrow.

Tuesday, January 24, 2006

Installing the x-ray gun

We input the x-ray gun this morning. Then I measured the resistance between the anode and ground, it's about 50G Ohms. Then I turn on the vacuum pump to evacuate the x-ray gun, I found the resistance decreased to 1M Ohms when the vacuum reached low than 1 Torr. But the resistance was back to 50G Ohms after the vacuum was under 10^-3 Torr. The cacuum was 2x10^-5 Torr when I left lab.

Monday, January 23, 2006

Cleaning the anode back epoxy bushing

We withdrew the X-ray gun even it had been taken out last month. I opened the back plate and removed the back insulator. It looked very dirty, and the measured resistance was less than 6M Ohms. Undoubtedly the back bushing is the key to make the x-ray anode trigger failure. It's incredible that the back bushing bacame black dirt after running a month. I think it may be caused by the surface grease when we installed it? By the way, the Viton O-ring 348 to sealthe epoxy block and aluminium plate looked coarse, maybe this caused gas leaking? When I removed the x-ray cathode bar, I found one of the binding nuts used two o-rings for sealing, maybe this carelessness caused the leakage?

We polished the epoxy bushing using the scouring pad and cleaned it with Kimberly wipers and ethanol. After all the resistance became larger than 50G Ohms.

Otherwise, the resistance between cathode and ground became 1.6M Ohms, it's very low. So we removed the cathode bars to clean them again, make sure the cathode is open to the ground.

Friday, January 20, 2006

Cleaning X-ray Anode Bushing

Mike and I measure the resistance between x-ray anode and ground using the megohm meter, it was about 5.8M Ohm as applied 1kV test voltage. It's definite that the front or back bushing was not good insulated. So I took out the anode bar to check the front bushing. I found the epoxy insulator looked very dirty, there was oil track on the surface. I think the silicon oil might leak inside the x-ray gun when we installed the anode bar. We cleaned and repolished the bushing, it looked more cleaning(shown as the top picture).

I measured the resistance, which was 50G Ohms. But the measured resistance became 5.8M Ohms after we inserted the anode bar into the x-ray gun. I will have to pull out the x-ray gun to check the back bushing next week.

Thursday, January 19, 2006

Problem for X-ray Anode

I opened the X-ray anode 4 caps oil tank, and took the X-ray Anode trigger signal sampling circiut board(shown as right picture). This voltage monitor plate is different from what I imaged. This plate is placed near the high voltage cable, the cable radiation will induce a small voltage through capacitors. This induced signal is used to measure the x-ray anode pulse. The circuit drawing is revised correspondingly, please check this link. As a result, we don't worry about the voltage divider as we thought yesterday.

I pulled out the 4 caps and checked them separately, we did not find anything wrong. Then Mike suggested to measure the high voltage cable, so we disconnected the cable from the x-ray anode bar and applied 10 kV to the cable. The measured current was less than 0.1 mA, which means the cable is good. Up to now, we only left one item was not detected - X-ray anode bar. Thus we directly connected the anode bar without 4 caps. The measured current increased up to 1 mA as we slowly tuned the high voltage to 10 kV. We concluded that there must be a fixed resistance between anode bar and ground, the value was estimated by ~10M Ohms. As we kept increasing voltage to 20 kV, the current measured up to 1.5 mA, then droped a little bit. The current became unstable, swung between 1.2 mA and 1.5 mA. There may be two reasons to cause this phenomenon: 1) the flashover was occured on the surface of conical epoxy bushing; 2) discharging between aonde pipe and cathode bar because of the anode pipe bending .

Wednesday, January 18, 2006

H.V. for thyratrons OK! Check X-ray Anode trigger independently

I pumped the oil from top tank and checked connects inside the oil tank carefully. I found where the problem was finally. Then main point is the voltage monitor connection. This voltage divider should be connected after the charging capacitors(black line), but it was put before them(red line, wrong connect). It's impossible to charger the caps, the voltage were drained down through resistors before charging the caps. I remembered that I installed the north side by myself after cleaning, and the south side was connected by other guys. I have not checked the wire connection after installing whole system. This gave me a lesson seriously, which is that I must check the circuit carefully anywhere and anyhow.

The system ran well in the afternoon except the x-ray anode randomly triggered. Mike suggested measuring the anode high voltage power supply independently. Firstly we only gave the power to 30kV HV power supply, in this case we don't worry about the low level trigger. When I increased the voltage slowly to 10kV, the current meter showed about 1~2 mA steadily. From the circuit drawing, the current should keep lower than 0.5 mA. We guessed the resistors might have problem. Then we disconnected the cable between transformer and 4 caps (shown as in the picture) and checked the front end circuit separately. This time we increased the voltage gradually to 30 kV, the current always was kept around 0.2 mA. Now it's pretty sure the breakdown was came from either caps, x-ray anode voltage divider or x-ray anode bar. The 4 caps are brand new, and just were replaced last month. They can be excluded. We will check the voltage monitor for x-ray anode tomorrow.

I also changed the O-ring for oil circulation system. I found the O-ring became crispy so that caused the oil leak from the seal slit. The O-ring model is 261, Kevin helped me to find it from the work shop.

Drawing list

I sent the 24 pieces of schematics not including mechanical drawing to Randy for scanning. The drawing list is shown below.

  13. C2271 H.V. CONTROL P.C. BOARD
  21. C2377 H.V. CONTROL P.C. BOARD

Tuesday, January 17, 2006

HV Power Supply

The resistance of the big transformer primary and secondary were measured, the value of the primary was zero, and the secondary 2k Ohm. Keith said that values are reasonable. I also checked the H.V. after disconnecting the cable to the thyratrons. The voltage obviously could be monitored. It means the HV power supply works very well. It must be somewhere not good after the HV power supply. I pumped the top oil tank to check the caps, chokes and transformers again. Because the wired phenomenon of no HV monitor was happened just after we cleaned the top tank oil and installed the caps last month.

Changed the dry compressed air cylinder for spark gap switch. Changed some damaged resistors.

Friday, January 13, 2006

High Voltage Board

We bypassed the SCR CD430440 to measure the transformer primary input voltage, the value was ~200V as we turn on the H.V. switch. But the HV meter was 0 V. Mike guessed the rectifiers may be broken, so we removed them for detecting. We applied a +1kV DC and -20 V DC to the rectifier one by one, we did not any one was damaged. We also checked the connect of 10 k Ohm resistors, and have not fond anything wrong. The resistor for current monitor also measured, the value is 70 Ohm. Till now we leave transform and voltage divider not be checked.

Thursday, January 12, 2006

Pulse Modulator H. V. Power Supply

I checked the X-ray Anode Control Box (MO5A) this morning. Mike suggested me to monitor the trigger signal of Anode H.V. Control PC Board (circuit drawing C-2302). We found the optically couple isolator phototransistor 6N136 was broken. After replaced the damaged chip, the X-ray anode can be triggered.

However, the 30 kV for thyratrons still could not be applied. Mike and I checked the low level circuit and did not find anything wrong. Then we firstly measured the voltage cross relay K4, it's ~205 V as the relay switched on. We were afraid that the SCR could not be controlled, so we monitored the voltage crossed SCR, it's ~200 V when the control panel H.V. switch was off. The value became to 0 V as the H.V. switch was on. It means the SCR and control system working correctly, but the H.V. meter show the voltage was 0 V. We will check transformer and rectifiers tomorrow.

Wednesday, January 11, 2006

No voltage supplied for Anode Hiptronics power supply

The repaired Lecroy oscilloscope was sent back this morning, we can use it to record the waveforms from now.

I received the new relays this afternoon, the circuit breaker was still kicked off after settling down them. Mike suggested to check the voltage phases. I found that I connected K3-1 and K3-3 incorrectly. The filament voltage became normal(~5.4 V DC, each side 2.7 V) after switching these two connectors. Then I tried to test the whole system, but this time there was no voltage spplied to Hiptronics High Voltage Power Supply. I measured the voltages on the MO5B board, the TB01-01 and TB01-03 both are zero, they are connected to MO5ATB03-01 and MO5ATB02-09 respectively(circuit drawing shown right). So I will check MO5A tomorrow.

News: Laser-driven proton

Recently, it has been observed that intense, collimated beams of multi-MeV protons were generated in the forward direction by intense short-pulse lasers irradiating thin solid foils at the Laboratoire pour l’Utilisation des Lasers Intenses(LULI) of CNRS. This discovery of an intense (~10^11 protons/shot with a maximum energy of ~25 MeV), pulsed (duration less than 1 ps), laminar source of protons has opened new and fascinating perspectives.

he production of such high current protons and ions beams have been made possible by short-pulse, ultra-high intensity lasers (laser pulse duration less than 1 ps, power density more than 10^18 W/cm^2 um^2) such as the one at LULI(see photo on the right showing the compressor and target chamber). There are several mechanisms that can lead to laser-acceleration of protons in the forward direction. In experiments performed at LULI, it has been unambiguously demonstrated that rear-surface acceleration is the predominant mechanism to get protons of high energy. In this mechanism, relativistic electrons generated at the laser-irradiated interface by the intense laser and propagated through the target form on the non-irradiated rear-surface of the target a dense electron plasma sheath. Its strong (~TV/m) electric field can ionize atoms and rapidly accelerates ions normal to the initially unperturbed surface. The accelerated protons stem either from hydrogen-containing contaminants (e.g. water vapour) that are present on all target surfaces or from prepared layers.

J. Fuchs, et al,Laser-driven proton scaling laws and new paths towards energy increase, Nature Physics 2, 48-54 (2006).
Laser Interaction group of LULI and Ecole polytechnique

Tuesday, January 10, 2006

Cleaned the Relays

The ordered relays have not been received until this afternoon, so I tried to clean the old relays by myself. Because the pin surfaces were burned very much, I have to file them so that pin thickness became less than before. In order to check the refurbished relays, I put them back to test, however the circuit breaker shut down once again. Then I disconnected the transformer output to purify the test condition, I found the relay K2 was close in the beginning. After 1 minutes time delay the relay K3 should be closed, but the circuit breaker was shut down while K3 was reacting. I estimated the primary problem is K3 caused circuit breakdown. The refinished relay K3 may not be used again.

Monday, January 09, 2006

History of Prometheus

I removed relays K2 and K3 from the MO4 circuit board. They looked very dirty, I disassmbled one and found some copper pins were burn into black. I am not sure these relays can be used again after cleaning the dirty pins, I think it's better to use new relays. Unfortunately I have not reveived them from Newark.

Randy wrote down a report about Prometheus, the followed is the part about its history. Thanks Randy!
History of Prometheus (by Randy Carlson)

Prometheus is one of two identical amplifiers delivered to both Los Alamos and UIC in the spring of 1988 by Beta Development Corporation under Gary Loda (now deceased). The device consists of two main discharge cavities (10cm x 10cm x 2m long) that share a common x-ray pre-ionizer that operates at 150 kV. Originally, each main discharge electrode was driven by a two-stage saturable core compressor with the main voltage step-up achieved by a 6.33:1 Stangenes autotransformer. The prime energy store is at 25 - 30 kV and switched to each discharge side (called North and South) by a North and South bank of six Thyratrons. This design architecture was chosen by the manufacturer as a way to achieve ~ 1 Hz or more repetition rate operation and results in a final fast rising voltage pulse of 150 ? 180 kV on the main discharge anodes. At Los Alamos and UIC the discharge cavities were utilized in a folded series configuration except that Los Alamos used a XeCl based gas mixture at 3 atm and UIC uses a KrF based gas mixture at about 1.6 to 1.7 atm.

Shortly after installation, two main changes were made to the x-ray pre-ionizer: (1) the smooth conical epoxy bushing that had evidence of flashover was machined with grooves, and (2) x-ray masks were utilized in an attempt to control the area for main energy disposition in the laser cavity. The first of these changes is present in the UIC amplifier but as for the second change. Another area of change was the replacement of both stages of saturable cores and one set (an array of 80 ceramic capacitor stacks) with two Maxwell Rail Gaps. Both sides of the UIC amplifier have the Maxwell rail gaps installed and other than annual maintenance are working in a flawless fashion. The Los Alamos system was salvaged in the mid-1990s.

Friday, January 06, 2006

Randy visited lab

Randy visited our lab today, he gave me many suggestions about high voltage pulsed power. I told him how the system run, what's main problem we have and gave him the documents and circuit drawings.

I wanted to show Randy the X-ray gun triggering, however the MO4 circuit breaker was shut down just after several minutes. I guessed the relay K2 or K3 was not good, because I have not changed them yet. I hope I can receive the parts next Monday.

Thursday, January 05, 2006

Wavelength Measurement of Coherent Innova 90

We ran the old Coherent INNOVA 90 Ar ion laser for new interferometer measurement in the new target chamber. We are not sure the output wavelength, so I used the CVI spectrometer to measure it.

Wavelength selection is dependent on the type of rear mirror holder. The INNOVA 90 can be operated with either of two types of rear mirror holders: a multiline mirror holder which produces operation in the all lines, or the Model 934 Wavelength Selector which produces operation in the single line.

For multiline operation, the multiline mirror holder should be used. When switching from single line operation to multiline operation, peak the output at 488.0 nm for argon system. The wavelength calibratiors located on the control panel should be set to read 488 nm for multiline argon operation.

Wednesday, January 04, 2006

Relay K3

I ran the system after tightening the charging resistor connector, unfortunately the high voltage still could not apply to the caps. The circuit breaker (CB1 as shown in the picture) was switched off after several minutes. Mike helped me to check the circuit drawing, we guessed the problem may be from the relays K2 and K3, so I measured the voltages crossing them. The relay K2 worked well. But the voltage cross resistor R1 was ~30V, it means the K3-1 connect is very bad.

I removed the relay K3, found the connect pin was very dirty and the connect surface was so rough. The shape of one pin was changed, so that it could not attached the wire connector well. The relay model is P25P42A22P1-120, the manufactory is Potter & Brumfield. I ordered it from Newark online.

Tuesday, January 03, 2006

Checked the system

I ran the Prometheus this afternoon, however the x-ray gun still was unstable. The new problem was the high voltage could not be applied to the rail gap. The filament switch was turn off after running several mintues, then I found the fuse was broken. I measured the Primary voltages, the north bank looked well, but there was no signal from south side.

I changed the fuse and the one 27 Ohm resistor for Thyratron at South bank.